OpenStudy (anonymous):

Determine the value of "r" so the line that passes through each pair of points has the given slope (m). (6,8), (r,-2), m = -3

jimthompson5910 (jim_thompson5910):

Hint: Solve (-2-8)/(r - 6) = -3 for r

OpenStudy (anonymous):

Ok so I have -10 / r-6 = -3. So r-6 has to equal one third of 10 for the fraction to work out to -3, so does that mean the answer is r = 9.333333333...?

jimthompson5910 (jim_thompson5910):

sry the website wouldn't load, but it seems to be working now

jimthompson5910 (jim_thompson5910):

yes r = 9.3333, but I would keep it in fraction form so r = 28/3

OpenStudy (anonymous):

Ok thanks!

jimthompson5910 (jim_thompson5910):

you're welcome